OpenStudy (anonymous):

Give me the steps not the answer please. How would i write log6 5 as a logarithm of base 4?

OpenStudy (amistre64):

well, write it out first log6 5 = log4 x swap these into lns by the change of base ln5 ln4 --- = ---- ln6 lnx then rearrange to solve for x

OpenStudy (anonymous):

ok, yea, i dont get it

OpenStudy (amistre64):

have you heard of the change of base ?

OpenStudy (anonymous):

no

OpenStudy (amistre64):

hmm, well then we can work it the long way i spose ... maybe

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

hmm, there is no pretty little answer to it is there .....

OpenStudy (anonymous):

no, this is confusing

OpenStudy (amistre64):

\[log_6(5)=log_4(x)\] \[\frac{ln5}{ln6}=\frac{lnx}{ln4}\] \[\Large e^\left(ln4\frac{ln5}{ln6}=lnx\right)\] \[\Large e^\left(ln4\frac{ln5}{ln6}\right)=x\] \[\Large e^\left(ln4\frac{ln5}{ln6}\right)=x\] which is another way of writing:\[\Large 4^{log_6(5)}\]

OpenStudy (anonymous):

this?

OpenStudy (amistre64):

well, that would work as a change of base yes .... i spose we aint got to use the lns

OpenStudy (amistre64):

5 on top tho

OpenStudy (anonymous):

thanks!

OpenStudy (amistre64):

yw